OpenStudy (anonymous):

Project Option 1 Designing a Park

OpenStudy (anonymous):

• City maps are often shown on a coordinate grid to make them easier to use. A central location can be set as the origin of the grid, and the regions of the grid can be framed by roads or other features to create closed shapes. For this option, you will design a new park by creating a colored-in, closed figure on a coordinate plane. Create a fun figure that will have everyone coming to your park. A diamond, a trapezoid—you name it! Then, you will state the inequalities that allow you to create the figure. You may use the link and directions below to help you create your shape or you may use any paint program, slideshow program, or graph the figure by hand, scan, and attach. Project Option 1 Designing a Park • City maps are often shown on a coordinate grid to make them easier to use. A central location can be set as the origin of the grid, and the regions of the grid can be framed by roads or other features to create closed shapes. For this option, you will design a new park by creating a colored-in, closed figure on a coordinate plane. Create a fun figure that will have everyone coming to your park. A diamond, a trapezoid—you name it! Then, you will state the inequalities that allow you to create the figure. You may use the link and directions below to help you create your shape or you may use any paint program, slideshow program, or graph the figure by hand, scan, and attach. For help with creating a shape in GeoGebra, select Geogebra directions to learn. You must submit the following to your instructor: 1. The graph: A coordinate plane with a colored in, closed figure. Each segment must be labeled. 2. The inequalities: Show the work for finding the inequality that defines each segment. 3. A point: Choose a point within the boundaries of the shape, and show algebraically that the point satisfies all of the inequalities. 4. Your reflection: Write a paragraph of three or more sentences as you reflect on this activity and your final product. You may use specific questions to guide your thoughts, if needed.

OpenStudy (anonymous):

@IMStuck @texaschic101

OpenStudy (anonymous):

hey tex

OpenStudy (anonymous):

know where IMstuck is?

OpenStudy (texaschic101):

who knows...lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

so, like i said, its a big project

OpenStudy (anonymous):

and i have no idea how to do it

OpenStudy (anonymous):

just sent a message to undeadknight so he may or may not show up

OpenStudy (anonymous):

so...what now

OpenStudy (anonymous):

texas should i just post this again later?

OpenStudy (texaschic101):

I have something.... y <= x - 6 y <= -x - 6 y >= x - 6 y >= -x - 6 when graphing these inequalities, the solution is in the shape of a diamond